OpenStudy (wolfboy):

Aysha has 8 pictures to hang over her couch. She wants to hang only 3 of them. Find how many ways she can choose the 3 pictures from the 8

OpenStudy (anonymous):

8p3

OpenStudy (anonymous):

8c3 if order doesn't matter.

OpenStudy (anonymous):

Are you familiar with Pascal's Triange?

OpenStudy (wolfboy):

uh not realy

OpenStudy (anonymous):

u kno wat 8p3 is?

OpenStudy (wolfboy):

not exactly

OpenStudy (anonymous):

I'm pretty sure this is a combination and not a permutation.

OpenStudy (anonymous):

yup that's a combination nd it's 8c3

OpenStudy (anonymous):

Can use either Pascal's Triangle or the formula with factorials. For n=8, probably better off using the formula.

OpenStudy (anonymous):

nCx = n!/(x!(n-x)!) I think that's it. I'm just pulling it from memory, so I might have it mixed up with the other one, but it looks right.

OpenStudy (wolfboy):

ok so how do we use this to make our equation. And sorry for the slowness i went surfing yesterday and i am still really tiered.

OpenStudy (anonymous):

A quick drawing of Pascal's Triangle shows it's 56, and 8!(3!(8-3)!) = the same.

OpenStudy (anonymous):

Do you know how to write out a factorial?

OpenStudy (wolfboy):

uh i probably do but dont remember can you show me again?

OpenStudy (anonymous):

While I do that, do an internet search for "Pascal's Triangle." - It's useful not only for combinations, but also for finding the coefficients of an expanded power of a binomial.

OpenStudy (anonymous):

n! = n*(n-1)*(n-2)*...*1 e.g. 8! = 8*7*6*5*4*3*2*1 dividing factorials is easy because so many common factors cancel out. e.g. 8!/3! = 8*7*6*5*4

OpenStudy (wolfboy):

Thanks for all of your help @CliffSedge